RR 10 breaths/min, I/E ratio is 1:2 Total Cycle Time:_ IT: ET: RR 12 breaths/min, VE ratio is 1:4 Total Cycle Time: IT: ET: RR 20 breaths/min, VE ratio is 1:1 Total Cycle Time: IT: ET: Set RR 15 breaths/min, set Vt 0.82L, flow rate 60 L/min TCT- IT= ET= I/E ratio = Set RR 13 breaths/min, set Vt 750 mL, flow rate 75 L/min VE = TCT IT= ET I/E ratio =
Answer #1

ANSWER

* RR 10 beat per breath, I/E ratio is 1:2.

        * Total cycle time (TCT)

                      Formula = 60 / RR

                        RR = 10 breath /min.

                         * Total cycle time = 60 / 10

                                                  = 6 sec.

          * TCT is equal to the sum of IT and ET.

                                    so   I/E ratio is 1:2

                                       IT = 2 sec.

                                       ET= 4 sec.

* RR 12 breath/min, I/E ratio 1:4.

                  * Total cycle time.

                         TCT = 60 / RR.

                                 = 60 / 12

                                 = 5 sec.

                  * TCT = 5 sec.

                      I/E Ratio = 1:4.

                          So

                              IT = 1 sec.

                              ET = 4sec.

* RR is 20 breath/min, I/E ratio is 1:1.

                       * TCT = 60/RR

                                   = 60/20

                                   = 3 sec.

                       * TCT = 3 sec.

                             I/E ratio = 1:1

                               So

                                   IT = 1.5 sec.

                                   ET = 1.5 sec.

* RR 15 breath/min, Vt 0.82 L, Flow rate 60L /min.

                     * Minute volume (VE).

                                   Formula

                                          VE = Tidal volume (Vt) * Respiratory rate (RR).

                                   Here

                                         Tidal volume = 0.82 L

                                   Respiratory rate = 15 breath /min.

                                                VE = 0.82 * 15

                                                      = 12.3 ml.

                  * TCT

                         TCT = 60 / RR

                                 = 60/ 15

                                 = 4 sec.

                  * IT

                      Formula = Tidal volume / flow rate * 60

                          IT = 0.82 /60 * 60

                              = 0.014 * 60

                              = 0.84 sec.

                 * ET

                         ET = TCT - IT

                              = 4 - 0.84

                              = 3.16 sec.

                 * I/E ratio

                             For I/E ratio divide both IT and ET with IT.

                               I/E ratio = 0.84 / 0.84 : 3.16 / 0.84

                                            = 1 : 3.8

* RR 13 breath /min,Vt 750 ml, Flow rate 75 l/min.

                    * Minute volume (VE)

                                VE = Vt * RR

                                Vt = 750 ml (0.75L).

                                RR = 13 breath /min.

                                 VE = 0.75 * 13

                                       = 9.75 ml.

                   * TCT

                          TCT = 60 /RR

                                  = 60/ 13

                                  = 4.6 sec.

                 * IT

                          IT = Vt / flow rate * 60

                              = 0.75 / 75 *60

                              = 0.01 * 60

                              = 0.6 sec.

                  * ET

                         ET = TCT - IT

                               = 4.6 - 0.6

                               = 4 sec.

               * I /E ratio

                          For I /E ratio divide IT and ET with IT.

                               I / E ratio = 0.6 /0.6 : 4 / 0.6

                                              = 1 : 6.67